OpenStudy (anonymous):

If coshx=5/3 and x>0, find the values for sinhx, tanhx and cschx

OpenStudy (anonymous):

I know it's hyperbolic, and there are formulas involved. Just not sure how to actually do the problem.

OpenStudy (tkhunny):

\[\cosh^{2}(x) - \sinh^{2}(x) = 1\] That should do for sinh(x)

OpenStudy (anonymous):

first take square cosh^2x= 25/9 for sinh^2x=1-cosh^2x as sinh^2x+cosh^2x=1 take sqreeroot to get sinhx now cschx=1/sinhx tanhx=sinhx/coshx

OpenStudy (anonymous):

why would i use that formula?

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

for sinhx

OpenStudy (anonymous):

so sinhx should = -4/3

OpenStudy (anonymous):

4i/3

OpenStudy (anonymous):

right,

OpenStudy (anonymous):

tan = 4i/5 and csc = 3/4i?